ISOMERIZATION, PERTURBATIONS, CALCULATIONS AND THE S_{1} STATE OF C_{2}H_{2}

\underline{J. H. BARABAN}, P. B. CHANGALA, J. R. P. BERK, R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139, USA; J. F. STANTON, Institute for Theoretical Chemistry, Departments of Chemistry and Biochemistry, The University of Texas at Austin, Austin, Texas 78712; A. J. MERER, Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan. Department of Chemistry, University of British Columbia, Vancouver, B.C., Canada V6T 1Z1.

Preliminary analysis of the energy region of the cis-trans isomerization transition state on the S_{1} surface of C_{2}H_{2} has revealed novel patterns and surprising perturbations, including unusually large (and high-order) anharmonicities, as well as K-staggerings of several vibrational levels. These effects complicate the analysis considerably, and require new models and calculations to account for and predict features of the observed spectra. The Ã − X̃ spectrum of acetylene has been studied both experimentally and theoretically for almost a century, and this cycle of unexpected phenomena eliciting innovative responses is found throughout its history. Especially in the last ten years, progress in understanding the S_{1} state rovibrational level structure and cis-trans isomerization has been accelerated by combining the information available from both ab initio computation and spectroscopic observations. The resulting dialogue has then frequently suggested fruitful avenues for further experiments and calculations. Current challenges and recent results in understanding the cis-trans isomerization transition state region will be discussed in this context.